OpenStudy (anonymous):

What would be the resulting molarity of a solution made by dissolving 45.7 grams of Mg(OH)2 in enough water to make a 1087-milliliter solution? Show all of the work needed to solve this problem.

OpenStudy (jfraser):

Use the molar mass to find moles of Mg(OH)2. Convert the volume of the solution to Liters.\[M = \frac{mol \space solute}{L \space solution}\] plug.

OpenStudy (anonymous):

moles Mg(OH)2 = 45.7 g /58.319 g/mol=0.784 M = 0.784 / 1.087 L=0.721